OpenStudy (anonymous):

Evaluate the integral: x(x^2-1)^4 dx

OpenStudy (anonymous):

\[\int\limits_{}^{}x(x^2-1)^4 =\int\limits_{}^{} u^4 1/2 du , u = x^2-1, = 1/10 (x^2-1)^5\]

OpenStudy (anonymous):

Thank you!